Ben is 12 years older than Lisa. The sum of their ages is 64. How old is Ben?
Answer by stanbon(75887)

Ben is 12 years older than Lisa.
The sum of their ages is 64.
How old is Ben?
---------------------------
B = L + 12
B + L = 64
-----------------------
Substitute for "B" and solve for "L"::
L + 12 + L = 64
------
2L = 52
------
L = 26 yrs (Lisa's age)
B = L+12 = 38 yrs (Ben's age)
========================
Cheers,
Stan H.